OpenStudy (anonymous):

Solve for x: log (x^3)=(log x)^2. Note, there are 2 solutions, A and B, where A

OpenStudy (freckles):

use power rule for log

OpenStudy (freckles):

log(x^n)=nlog(x)

OpenStudy (freckles):

\[0=(\log(x))^2-3\log(x)\] then factor the expression on the right and solve for log(x)

OpenStudy (freckles):

then solve for x